Write the equation of the line, in point-slope form and slope-intercept form, that contains the points (-2, -1 ) and (2, 1 ). Use the box provided or the upload option to submit all of your calculations and final answers.

Answer by Alan3354(69443)

Write the equation of the line, in point-slope form and slope-intercept form, that contains the points (-2,-1) and (2,1).
------------
By inspection, it's symmetrical about the Origin.
--> y = x/2
==================
Find the slope.
slope m = diffy/diffx = 2/4 = 1/2
----
y-y1 = m*(x-x1)
y-1 = (1/2)*(x-2)
y = (1/2)x
--------------
=============
PS Don't put spaces in the points.